OpenStudy (jziggy):

In a process that manufactures bearings, 90% of the bearings meet a thickness specification. A shipment contains 500 bearings. A shipment is acceptable if at least 440 of the 500 bearings meet the specification. Assume that each shipment contains a random sample of bearings. What is the probability that a given shipment is acceptable?

OpenStudy (jziggy):

What I've tried so far is a normal approximation to the binomial X ~ Bin(500 , 0.9) Which goes to x ~ N(450 , 6.71^2) Then I found P(x>440) = 1 - p(x<440) = 1 - Phi(-10/6.71) = Phi(1.49) Which from the z-table is .9319 According to the back of the book, the answer is supposed to be 0.9418. What did I do wrong?

OpenStudy (kropot72):

In this question the continuous Normal distribution is being used to approximate a Discrete distribution. Therefore the Correction for Continuity must be applied, giving: \[\large P(X \ge 440)=1-\phi(\frac{-10.5}{6.71})=you\ can\ calculate\]